# ZEUS: Zero-shot Embeddings for Unsupervised Separation of Tabular Data

### v1

@@ 1 Introduction
Clustering remains a fundamental yet challenging task in unsupervised learning. It is particularly hard for tabular data, which inherently lacks the structured spatial or semantic properties of images or texts. Unlike for image clustering, where intrinsic visual similarities can guide cluster formation, defining meaningful similarities in tabular data is highly dataset-specific, complicating the generalization of clustering methods across diverse applications. Recent developments leveraging deep learning have demonstrated promise in generating richer representations for clustering tasks. However, these methods frequently suffer from instability due to their sensitivity to hyperparameter selection, a challenge exacerbated by the absence of supervised signals to guide optimization. Consequently, practitioners working with tabular data often resort to simpler, classical algorithms like k-means, despite their limited capacity for capturing complex underlying data structures, simply to avoid extensive manual tuning.
@@
We address these challenges with ZEUS – a zero-shot transformer-based model for embedding new tabular datasets in a form convenient for unsupervised separation (=clustering), without the need for additional fine-tuning. Given a new dataset, ZEUS returns its transformed representation, where clusters can be easily discovered using simple methods, like k-means. Since it works as a zero-shot learner, it significantly reduces hyperparameter tuning complexity and the computation time, enabling effective clustering in seconds. It is a plug-and-play solution that requires no fine-tuning and runs in a single forward pass for new datasets.
@@
Inspiration for ZEUS stems from the Prior-data Fitted Networks (PFNs) tabpfn , a recently introduced framework highlighting the potential of in-context learning for tabular data. PFNs operate by feeding a transformer contexts representing complete tasks, i.e., training data along with query samples for which the model predicts labels. Despite their impressive performance, they are limited to supervised problems. ZEUS extends this basic idea for unsupervised tasks by addressing the two key challenges of how to: (1) generate prior synthetic data with clear but non-trivial clustering structures for pre-training; and (2) encode prior clustering knowledge for new unlabeled datasets. Additionally, in contrast to TabPFN, ZEUS is a zero-shot model and during inference does not rely on any context labels.
@@
Unlike traditional methods that optimize arbitrary heuristics (e.g., DEC), ZEUS approaches clustering by learning how to invert data generation processes. In particular, it is pre-trained on synthetic datasets to learn how to infer cluster assignments. The datasets have known latent structures, which enables supervised guidance for the training, and by generating diverse datasets we supply it with the prior knowledge about what can constitute a possible clustering structure, enabling effective generalization to new real-world datasets during inference. Figure 1 illustrates the key concepts of the method while Figure 2 presents representations generated by ZEUS.
@@
Experiments demonstrate that ZEUS consistently matches or surpasses both classical clustering algorithms and recent deep learning-based approaches, establishing it as a powerful and practical tool for the unsupervised analysis of tabular data. Beyond its strong empirical performance, we also provide theoretical justification by showing that ZEUS fits the framework of Prior-data Fitted Networks, thereby reinforcing its theoretical soundness (see Section 2.3).
@@
We supplement the paper with an appendix that includes background on prior-data fitted networks, details of the synthetic data generation process, and additional details of the experimental results. The code used in this paper is available at https://github.com/gmum/zeus
@@ 2.1 ZEUS method
Traditional machine learning models $f_{\theta}$ are trained by optimizing parameters $\theta$ to solve a specific fixed task, such as classification or clustering, over a fixed dataset. On the other hand, in-context learning (ICL) 10.5555/3600270.3602487 ; lee2023supervised ; han-etal-2023-understanding enables pre-trained models $f_{\theta^{\ast}}$ to adapt to new tasks or datasets without ever updating the pre-trained parameters $\theta^{*}$. Similarly, zero-shot learning allows models to address problems they were never explicitly trained on, without requiring any examples of the target task. Unlike ICL, which adapts through demonstration examples, zero-shot approaches rely entirely on knowledge encoded during pre-training. The pre-training of zero-shot models typically involves massive and diverse datasets that encourage the model to learn rich, generalizable representations of underlying patterns and relationships. This usually is done using transformers, as they both can process set-valued inputs (e.g., accept a whole dataset at once) and model complex dependencies.
@@
For specifying $p_{y}(x)$ we draw an inspiration from Gaussian Mixture Models (GMMs) dempster1977maximum ; hansen1982large , where each cluster $k$ is represented by a Gaussian distribution centered at ${c}_{k}$, and data points are probabilistically assigned to clusters through probabilities $p_{k}(x_{i})$, based on their proximity to these centroids. Consequently, alongside the encoded representations, we need to define a set of $K$ cluster centroids (e.g., prototypical representations) $\{{c}_{k}\}_{k=1}^{K}$, which reside in the same representation space $S$. We then jointly optimize $\mathcal{L}$ w.r.t $\theta$ and $\{{c}_{k}\}$ such that the representations ${z}(x_{i})$ become structured in a way undisclosing underlying probabilistic structures in datasets.
We aim at higher probabilities $p$ for points closer these centroids. To achieve this, we first define scores $\alpha_{k}(x_{i})=-\|{z}(x_{i})-{c}_{k}\|^{2}$ to quantify the compatibility between representations and centroids ${c}_{k}$, and then transform them into cluster membership probabilities using softmax, yielding soft cluster assignments $p_{k}(x_{i})=\frac{\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\exp(\alpha_{j}(x_{i}))}.$ Note that the ground-truth cluster assignments $y_{i}$ are known during pre-training, and hence, the maximum likelihood centroids can be straightforward estimated as $\hat{c}_{k}=\frac{1}{N_{k}}\sum_{\{i:y_{i}=k\}}{z}(x_{i}),$ where $N_{k}$ denotes the number of data points assigned to the $k$-th cluster. Following this formulation, the optimization needs to performed only w.r.t the parameters $\theta$ of the neural network $f_{\theta}$, as the centroids are already specified by them.
@@
The structural relations learned during pre-training are leveraged during inference when the representations ${z}$ produced by the model $f_{\theta^{\ast}}$ are used to create a predictive distribution for new, previously unseen inputs $\mathbf{x}^{*}$ as $\hat{\mathbf{p}}^{*}:=p\bigl{(}\mathbf{y}^{*}\mid f_{\theta^{\ast}}(\mathbf{x}^{*})\bigr{)}.$ At this point, the labels $\mathbf{y}^{*}$ are unknown and therefore, cannot be used to estimate the centroiods nor the priors. Hence, to structure the obtained representation into clusters, we use a traditional learning algorithm, for example, similar to GMMs relying on Expectation-Maximization or simply k-means.
@@
Although theoretically $\mathcal{L}_{prob}$ is sufficient to build a clustering representation, we experimentally verified that the introducing additional regularizers further improves the predictions in the inference (see Section 3.4).
First, we aim to ensure that the representations ${z}(x_{i})$ associated with a particular cluster are compactly distributed around their corresponding centroid to enhance intra-cluster cohesion. A point concentration regularizer that explicitly minimizes the distance between representations and the centroids, analogous to the k-means objective of minimizing within-cluster sum of squares, we define as
@@ 2.2 Prior data for pre-training ZEUS
The key component of ZEUS is the data-generating prior $p(\mathcal{D})$. As it primarily affects the generalization ability of the pre-trained model $f_{\theta^{*}}$, it must be designed to cover diverse data distributions and various cluster configurations. In particular, we construct it as a latent variable model (LVM), assuming that each dataset $\mathcal{D}$ is sampled from a $K$-component mixture of distributions. Formally, the probability of a data point $x\in\mathcal{D}$ under this model is $p(x)=\sum_{k=1}^{K}p(y=k)\,p(x\mid y=k),$ where $p(y)$ is a categorical distribution, and $p(x\mid y=k)$ represents the $k$-th (continuous) component. Although the categories $y$ remain latent for real datasets, they are known during synthetic data generation and can serve as ground-truth labels – a property integral to pre-training ZEUS.
The number of categories $K$ we sample uniformly between 2 and 10, and the observations $x$ are sampled from multivariate Gaussian distributions. To control the complexity of datasets, we introduce a constraint that provides a sufficient separation between each pair of components. For each component, we generate between $50$ and $800$ samples. Since real data rarely consist of Gaussian clusters, we additionally transform the data points from each category using randomized ResNet-like neural networks to produce more realistic cluster shapes. We selected ResNets due to their properties, as they can define invertible transformations  (behrmann2019invertibleresidualnetworks, ), and this ensures that clustering structures will be preserved in outputs. Finally, we append a certain fraction of categorical features in one-hot encoding scheme to selected datasets, which is followed by an optional PCA reduction to keep the data dimension at the requested maximum level. Complete details of the prior-data generating process can be found in Appendix B.
@@ 2.3 Relation to Bayesian learning and Prior-Data Fitted Networks
[…]
ZEUS instead of directly outputting $q_{\theta}(y|x,\mathcal{D}_{ctx})$, maps inputs $x$ to latent representation vectors ${z}(x)$. The probabilistic assignments $p_{y}(x)$ are then constructed from these vectors according to Eq. 3. This equation specifies a PPD for inference, but also defines a probability mass function for training:
@@
As explained above our clustering-based extension is theoretically sound, nonetheless, the assumption could limit representational expressivity compared to the more flexible transformers employed for the original PFNs tabpfn . On the other hand, the enforced structures shall be more appropriate for unsupervised tasks.
@@ 3 Experiments
This section provides an experimental evaluation of the clustering performance of our method.
@@ 3.1 Experimental setup
Model architecture: ZEUS relies on a transformer architecture similar to TabPFN (tabpfn, ). It consists of 12 attention blocks, each with 6 heads and a token dimension of 512, employing GeLU as the activation function. Following the TabPFN approach, each data point is linearly embedded prior to the transformer component and then treated as a token.
Pre-training: In the pre-training phase, we sample datasets from the mixture of Gaussians and transformed mixtures in equal proportions (1:1), generating 1000 unique dataset batch samples for each epoch. For training, we employ the Adam optimizer along with a cosine learning rate scheduler with warm-up, using a learning rate of 2e-5. The plot illustrating model improvements during the pre-training process is available in Appendix C.
Inference: During inference, preprocessing of each dataset involves standardizing numerical features, followed by scaling them to the range $[-1,1]$, whereas categorical features are transformed using one-hot encoding. The input size of our model is fixed to 30. For datasets with lower dimensionality, we pad the missing positions with zeros, while for higher-dimensional datasets, we reduce the number of input features via Principal Component Analysis (PCA). Unless stated otherwise, at inference we use k-means applied to the normalized (=scaled to $[-1,1]$) transformer output in order to obtain clusters from our learned representation.
Datasets: For evaluation, we consider three groups of datasets: real datasets from OpenML (Real), synthetic mixtures of Gaussians (Syn. Gauss.), and synthetic mixtures of Gaussians transformed by ResNet-like neural networks (Syn. Transf.). Both types of synthetic datasets are augmented with categorical variables. The process of generating synthetic datasets is described in Section 2.2.
@@
Baselines: We compare ZEUS against a wide spectrum of state-of-the-art clustering methods used for tabular data. It includes $k$-means (KM), Gaussian Mixture Model (GMM), and deep-learning methods based on autoencoder (AE) architectures, including DEC xie2016unsuperviseddeepembeddingclustering , IDEC guo_improved_2017 , and G-CEALS rabbani2025deep . We additionally consider also $k$-means and GMM applied in the autoencoder latent space (respectively referred to as AE-KM and AE-GMM) to compare the clustering quality of our representation against the one obtained from the autoencoder.
@@
Reporting: In the main text, we report only the results aggregated for each group of datasets (Real, Syn. Gauss, Syn. Transf.), while detailed results for individual datasets can be found in Appendix F. For the reader’s convenience, we bold the best results and underline the second-best ones.

### v2

@@ 1 Introduction
Clustering remains a fundamental yet challenging task in unsupervised learning. It is particularly hard for tabular data, which inherently lacks the structured spatial or semantic properties of images or texts. Unlike for image clustering, where intrinsic visual similarities can guide cluster formation, defining meaningful similarities in tabular data is highly dataset-specific, complicating the generalization of clustering methods across diverse applications. Recent developments leveraging deep learning have demonstrated promise in generating richer representations for clustering tasks. However, these methods frequently suffer from instability due to their sensitivity to hyperparameter selection, a challenge exacerbated by the absence of supervised signals to guide optimization. Consequently, practitioners working with tabular data often resort to simpler, classical algorithms like k-means macqueen1967some , despite their limited capacity for capturing complex underlying data structures, simply to avoid extensive manual tuning.
@@
We address these challenges with ZEUS – a zero-shot transformer-based model for embedding new tabular datasets in a form convenient for unsupervised separation (=clustering), without the need for additional fine-tuning. Given a new dataset, ZEUS returns its transformed representation, where clusters can be easily discovered using simple methods, like k-means. Since it works as a zero-shot learner, it significantly reduces hyperparameter tuning complexity and computation time, enabling effective clustering in seconds. It is a plug-and-play solution that requires no fine-tuning and runs in a single forward pass for new datasets.
@@
Inspiration for ZEUS stems from the Prior-data Fitted Networks (PFNs) tabpfn , a recently introduced framework highlighting the potential of in-context learning for tabular data. PFNs operate by feeding a transformer transformer contexts representing complete tasks, i.e., training data along with query samples for which the model predicts labels. Despite their impressive performance, they are limited to supervised problems. ZEUS extends this basic idea for unsupervised tasks by addressing the two key challenges of how to: (1) generate prior synthetic data with clear but non-trivial clustering structures for pre-training; and (2) encode prior clustering knowledge for new unlabeled datasets. Additionally, in contrast to TabPFN, ZEUS is a zero-shot model and during inference does not rely on any context labels.
@@
Unlike traditional methods that optimize arbitrary heuristics (e.g., DEC xie2016unsuperviseddeepembeddingclustering ), ZEUS approaches clustering by learning how to invert data generation processes. In particular, it is pre-trained on synthetic datasets to learn how to infer cluster assignments. The datasets have known latent structures, which enables supervised guidance for the training, and by generating diverse datasets we supply it with the prior knowledge about what can constitute a possible clustering structure, enabling effective generalization to new real-world datasets during inference. Figure 1 illustrates the key concepts of the method while Figure˜2 presents representations generated by ZEUS.
@@
Experiments demonstrate that ZEUS consistently matches or surpasses both classical clustering algorithms and recent deep learning-based approaches, establishing it as a powerful and practical tool for the unsupervised analysis of tabular data. Beyond its strong empirical performance, we also provide theoretical justification by showing that ZEUS fits the framework of Prior-data Fitted Networks, thereby reinforcing its theoretical soundness (see Section˜2.3).
@@
We supplement the paper with an appendix that includes background on prior-data fitted networks, specifics of the synthetic data generation process, details of the experimental results, and additional experimental studies. The code used in this paper is available at https://github.com/gmum/zeus.
@@ 2.1 ZEUS method
Traditional machine learning models $f_{\theta}$ are trained by optimizing parameters $\theta$ to solve a specific fixed task, such as classification or clustering, over a fixed dataset. On the other hand, in-context learning (ICL) garg2023transformerslearnincontextcase ; lee2023supervised ; han-etal-2023-understanding enables pre-trained models $f_{\theta^{\ast}}$ to adapt to new tasks or datasets without ever updating the pre-trained parameters $\theta^{*}$. Similarly, zero-shot learning allows models to address problems they were never explicitly trained on, without requiring any examples of the target task. Unlike ICL, which adapts through demonstration examples, zero-shot approaches rely entirely on knowledge encoded during pre-training. The pre-training of zero-shot models typically involves massive and diverse datasets that encourage the model to learn rich, generalizable representations of underlying patterns and relationships. This is usually done using transformers, as they both can process set-valued inputs (e.g., accept a whole dataset at once) and model complex dependencies.
@@
For specifying $p_{y}(x)$, we draw inspiration from Gaussian Mixture Models (GMMs) dempster1977maximum ; hansen1982large , where each cluster $k$ is represented by a Gaussian distribution centered at ${c}_{k}$, and data points are probabilistically assigned to clusters through probabilities $p_{k}(x_{i})$, based on their proximity to these centroids. Consequently, alongside the encoded representations, we need to define a set of $K$ cluster centroids (e.g., prototypical representations) $\{{c}_{k}\}_{k=1}^{K}$, which reside in the same representation space $S$. We then jointly optimize $\mathcal{L}$ w.r.t $\theta$ and $\{{c}_{k}\}$ such that the representations ${z}(x_{i})$ become structured in a way undisclosing underlying probabilistic structures in datasets.
We aim for higher probabilities $p$ for points closer to these centroids. To achieve this, we first define scores $\alpha_{k}(x_{i})=-\|{z}(x_{i})-{c}_{k}\|^{2}$ to quantify the compatibility between representations and centroids ${c}_{k}$, and then transform them into cluster membership probabilities using softmax, yielding soft cluster assignments $p_{k}(x_{i})=\frac{\exp(\alpha_{k}(x_{i}))}{\sum_{j=1}^{K}\exp(\alpha_{j}(x_{i}))}.$ Note that the ground-truth cluster assignments $y_{i}$ are known during pre-training, and hence, the maximum likelihood centroids can be straightforwardly estimated as $\hat{c}_{k}=\frac{1}{N_{k}}\sum_{\{i:y_{i}=k\}}{z}(x_{i}),$ where $N_{k}$ denotes the number of data points assigned to the $k$-th cluster. Following this formulation, the optimization needs to be performed only w.r.t the parameters $\theta$ of the neural network $f_{\theta}$, as the centroids are already specified by them.
@@
The structural relations learned during pre-training are leveraged during inference when the representations z{z} produced by the model $f_{\theta^{\ast}}$ are used to create a predictive distribution for new, previously unseen inputs $\mathbf{x}^{*}$ as $\hat{\mathbf{p}}^{*}:=p\bigl(\mathbf{y}^{*}\mid f_{\theta^{\ast}}(\mathbf{x}^{*})\bigr).$ At this point, the labels $\mathbf{y}^{*}$ are unknown and therefore cannot be used to estimate the centroids nor the priors. Hence, to structure the obtained representation into clusters, we use a traditional learning algorithm, for example, similar to GMMs relying on Expectation-Maximization or simply k-means.
@@
Although theoretically $\mathcal{L}_{prob}$ is sufficient to build a clustering representation, we experimentally verified that the introduction of additional regularizers further improves the predictions in the inference (see Section˜3.4).
First, we aim to ensure that the representations ${z}(x_{i})$ associated with a particular cluster are compactly distributed around their corresponding centroid to enhance intra-cluster cohesion. We define a point concentration regularizer that explicitly minimizes the distance between representations and the centroids, analogous to the k-means objective of minimizing within-cluster sum of squares:
@@ 2.2 Prior data for pre-training ZEUS
The key component of ZEUS is the data-generating prior $p(\mathcal{D})$. As it primarily affects the generalization ability of the pre-trained model $f_{\theta^{*}}$, it must be designed to cover diverse data distributions and various cluster configurations. In particular, we construct it as a latent variable model (LVM), assuming that each dataset $\mathcal{D}$ is sampled from a $K$-component mixture of distributions. Formally, the probability of a data point $x\in\mathcal{D}$ under this model is
$p(x)=\sum_{k=1}^{K}p(y=k)\,p(x\mid y=k),$
where $p(y)$ is a categorical distribution, and $p(x\mid y=k)$ represents the $k$-th (continuous) component. Although the categories $y$ remain latent for real datasets, they are known during synthetic data generation and can serve as ground-truth labels – a property integral to pre-training ZEUS. This procedure, coupled with Theorem 1 (in Appendix) ensures that a sufficiently expressive model trained on synthetic datasets can handle real datasets drawn from arbitrary mixtures of distributions.
The number of categories $K$ we sample uniformly between 2 and 10, and the observations $x$ are sampled from multivariate Gaussian distributions. To control the complexity of datasets, we introduce a constraint (Eq. 7 in Appendix) that ensures a sufficient separation between each pair of components. For each component, we generate between $50$ and $800$ samples. Since real data rarely consist of Gaussian clusters, we additionally transform the data points from each category using randomized ResNet-like neural networks to produce more realistic cluster shapes. We selected ResNets due to their properties, as they can define invertible transformations  (behrmann2019invertibleresidualnetworks, ), and this ensures that clustering structures will be preserved in outputs. Finally, we append a certain fraction of categorical features in one-hot encoding scheme to selected datasets, which is followed by an optional PCA reduction to keep the data dimension at the requested maximum level. Complete details of the prior-data generating process can be found in Appendix B.2. Theoretical justification of this procedure is presented in Appendix B.1.
@@ 2.3 Relation to Bayesian learning and Prior-Data Fitted Networks
Given a prior distribution $p(\mathcal{D})$ over (synthetic) datasets $\mathcal{D}$, a PFN parameterized by $\theta$ is trained to minimize the negative log-likelihood of predicting held-out labels within datasets sampled from this prior. The associated loss function is defined as: $\mathcal{L}_{\textit{PFN}}(\theta)=\mathbb{E}_{\mathcal{D}_{ctx}\cup\{(x,y)\}\sim p(\mathcal{D})}[-\log q_{\theta}(y|x,\mathcal{D}_{ctx})].$ Minimizing the Prior-Data Negative Log-Likelihood is then equivalent to minimizing the expected Kullback–Leibler divergence between the network’s predictive distribution and the true Posterior Predictive Distribution (PPD) $p(y|x,\mathcal{D}_{ctx})=\int_{\Phi}p(y|x,\phi)p(\mathcal{D}_{ctx}|\phi)p(\phi)d\phi$ (see Corollary 1.1 in bayes_in_context ). A pre-trained PFN approximates this integral implicitly, yielding a distribution $q_{\theta}(y|x,\mathcal{D}_{ctx})$ directly from forward propagation of the network. In Appendix˜A, a more detailed explanation of PFNs was provided.
ZEUS instead of directly outputting $q_{\theta}(y|x,\mathcal{D}_{ctx})$, maps inputs $x$ to latent representation vectors z(x){z}(x). The probabilistic assignments $p_{y}(x)$ are then constructed from these vectors according to Eq. 3. This equation specifies a PPD for inference, but also defines a probability mass function for training:
@@
As explained above, our clustering-based extension is theoretically sound, nonetheless, the assumption could limit representational expressivity compared to the more flexible transformers employed for the original PFNs tabpfn . On the other hand, the enforced structures shall be more appropriate for unsupervised tasks.
@@ 3 Experiments
This section presents an experimental evaluation of the clustering performance of our method. Due to space constraints, further results are provided in Appendix G.
@@ 3.1 Experimental setup
Model architecture: ZEUS relies on a transformer architecture similar to TabPFN (tabpfn, ). It consists of 12 attention blocks, each with 6 heads and a token dimension of 512, with GeLU activation employed. Following the TabPFN design, each data point is first embedded using a linear transformation and then passed as a token to the transformer. For the unsupervised setting, no label embeddings are ever created or presented to the model. Similarly, for the zero-shot case, no query set is used either, and consequently, attention is computed solely over the support set. Finally, unlike TabPFN, we omit any additional MLP decoder after the transformer.
Pre-training: In the pre-training phase, we sample datasets from the mixture of Gaussians and transformed mixtures in equal proportions (1:1), generating 1000 unique dataset batch samples for each epoch. For training, we employ the Adam optimizer along with a cosine learning rate scheduler with warm-up, using a learning rate of 2e-5. The plot illustrating model improvements during the pre-training process is available in Appendix˜C.
Inference: During inference, preprocessing of each dataset involves standardizing numerical features, followed by scaling them to the range [1,1][-1,1], whereas categorical features are transformed using one-hot encoding. The input size of our model is fixed to 30. For datasets with lower dimensionality, we pad the missing positions with zeros, while for higher-dimensional datasets, we reduce the number of input features via Principal Component Analysis (PCA). Unless stated otherwise, at inference we use k-means applied to the normalized (=scaled to [1,1][-1,1]) transformer output in order to obtain clusters from our learned representation.
Datasets: For evaluation, we consider three groups of datasets: real datasets from OpenML OpenML2025 (Real), synthetic mixtures of Gaussians (Syn. Gauss.), and synthetic mixtures of Gaussians transformed by ResNet-like neural networks (Syn. Transf.). Both types of synthetic datasets are augmented with categorical variables. The process of generating synthetic datasets is described in Section˜2.2.
@@
Baselines: We compare ZEUS against a wide spectrum of state-of-the-art clustering methods used for tabular data. It includes $k$-means (KM), Gaussian Mixture Model (GMM), and deep-learning methods based on autoencoder (AE) architectures, including DEC xie2016unsuperviseddeepembeddingclustering , IDEC guo_improved_2017 , IDC idc , and G-CEALS rabbani2025deep . We additionally consider $k$-means and GMM performed in the autoencoder latent space (respectively referred to as AE-KM and AE-GMM), along with $k$-means applied to representations obtained from TabPFN-Unsupervised hollmann2025accurate (TabPFN) and SCARF scarf , to assess the clustering quality of different feature representations.
@@
Reporting: In the main text, we report only the results aggregated for each group of datasets (Real, Syn. Gauss, Syn. Transf.), while detailed results for individual datasets can be found in Appendix˜F. For the reader’s convenience, we bold the best results and underline the second-best ones.